QUESTION 11 4 As the CFO of a company, you consider the following three projects: NPV 60.5K 66.6K 42.0K Project Investment 50K 60K 40K Unfortunately, the company has only S100,000 available for investment. Which projects do you choose based on weighted average profitability index (WAPI):? Proiect A alone Project A + Project B Project A + Project C Project B + Project C D.
Answer #1

We can calculate the profitability index of each project:

Profitability index (PI)=(Investment+NPV)/Investment:

Project NPV Inv PI
A 60.5 50 2.21
B 66.6 60 2.11
C 42 40 2.05

Hence we shall select A and B but total outlay shall exceed:

Hence if A and C selected, NPV=60.5+42=102.50

if B and C selected, NPV=66.6+42=108.6

Project B and Project C shall be selected.

Hence option D is correct